# Asymptotically Optimal Lazy Lifelong Sampling-based Algorithm for Efficient Motion Planning in Dynamic Environments

### v2

@@ IV-A Static Planning Problems
(a) Scene 1
(b) Scene 2
(c) Scene 3
Figure 3: The x-y plane of the obstacle configurations (represented by black regions) of the dynamic planning problems. The red dot and the black dot denotes the start and the goal configuration, respectively.
The simulation results are displayed in Fig.2, showcasing the medium solution cost plotted against run time. LLPT emerges as the top performer among all algorithms, excelling in both the efficiency of initial solution discovery and the convergence rate towards the optimal solution. These results demonstrate the efficacy of lazily delaying edge evaluations in solving motion planning problems. It’s worth noting that LLPT and RRTX outperform informed sampling-based methods (Informed RRT and BIT) in scenarios depicted in Fig.1(a), where the solution path is twisted. In such cases, the informed set comprises a considerable number of nodes, and the informed sampling strategies underperform. However, for the problems depicted in Fig.1(b), BIT displays superior performance over RRTX in $\mathbb{R}^{2}$, and marginally outperforms both RRTX and LLPT in $\mathbb{R}^{3}$ as run time increases. These findings suggest a potential avenue for performance enhancement by integrating LLPT with informed sampling strategies for scenarios with simple obstacle structures.
IV-B Dynamic Planning Problems
We compared the replanning efficiency of LLPT with RRTX and RT-RRT in solving dynamic path planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{3}$. During planning, the obstacle configurations change from Scene 1 to Scene 3 (Fig.3). The start and goal are located at $[15,1,15,...,15]^{T}$ and $[15,15,15,...,15]^{T}$, respectively. To ensure a fair comparison, all planners assume that the problem domain is free and initialize by generating an RRG $\mathcal{G}$ that approximates the problem domain and surely contains a solution for all cases. Specifically, $\mathcal{G}$ has 3000 for $\mathbb{R}^{2}$ and 9000 nodes for $\mathbb{R}^{3}$. Each planner then undergoes three consecutive planning iterations on $\mathcal{G}$ to adapt their solutions to the changing environment. Note that the environmental change interval is more significant than the time required for replanning to ensure that the planners have enough time to optimize their solutions. Each planner was run 20 times with different pseudo-random seeds.
The average number of edge evaluations, the average number of node expansions, the average time cost, and the average solution cost for initialization and each replanning episode are tabulated in Table I. In this paper, the number of node expansions refers to the total number of nodes that are popped from the priority queue during the replanning procedure. The results demonstrate that LLPT is able to replan successfully in all cases with significantly fewer edge evaluations while maintaining solution quality, showcasing a marked overall efficiency improvement. It’s worth noting that LLPT avoids unnecessary edge evaluations when updating its solution to Scene 1, unlike RRTX and RT-RRT, which perform iterative evaluations across the entire graph even when obstacle movement does not impact the solution.
Moreover, the impact of the evaluation parameter $\alpha$ on LLPT performance is explored in Table 1, revealing its influence on edge evaluations and node expansions. The paper suggests careful selection of $\alpha$ to optimize LLPT performance for specific problems, although the methodology for selecting the ideal value is beyond the paper’s scope.
Figure 4: LLPT planning results to find the shortest path in a dynamic environment. The purple lines are the evaluated edges. The blue lines represent the edges belonging to the search tree. The red line represents the solution path.
TABLE I: The average number of edge evaluations, the average number of vertex expansions, the average replanning time, and the average solution cost for different planners over three consecutive planning in the dynamic environment shown in Fig.3. The absence of a solution was considered a cost of 200.
$\boldsymbol{\mathbb{R}^{2}}$ LLPT RRTX RTRRT
$\alpha$=1 $\alpha$=3 $\alpha$=100
Initialize
# Eval. 10 10 10 119340 119340
# Expan. 799 799 799 11961 5728
Time (s) 47.5 47.5 47.5 67.7 67.0
Cost 14.1 14.1 14.1 14.1 14.1
1st Replan
# Eval. 10 10 10 119340 119340
# Expan. 0 0 0 304 2677
Time (s) 0.001 0.001 0.001 11.5 13.6
Cost 14.1 14.1 14.1 14.1 14.1
2nd Replan
# Eval. 205 341 415 119340 119340
# Expan. 3846 3530 3294 2104 2693
Time (s) 2.1 1.81 1.74 12.9 14.7
Cost 23.3 23.3 23.3 23.3 23.3
3rd Replan
# Eval. 562 852 974 119340 119340
# Expan. 44817 42799 38216 151 2660
Time (s) 14.4 13.5 12.8 12.8 14.5
Cost 51.8 51.8 51.8 200 51.8
Total Time (s) 64.0 62.8 62.0 104.9 109.8
$\boldsymbol{\mathbb{R}^{3}}$
Initialize
# Eval. 6 6 6 303122 303122
# Expan. 1486 1486 1486 20144 541619
Time (s) 430.0 430.0 430.0 516.7 521.4
Cost 14.8 14.8 14.8 14.8 14.8
1st Replan
# Eval. 6 6 6 303122 303122
# Expan. 0 0 0 823 8067
Time (s) 0.001 0.001 0.001 46.0 48.9
Cost 14.8 14.8 14.8 14.8 14.8
2nd Replan
# Eval. 475 798 839 303122 303122
# Expan. 6000 5766 5724 5914 8072
Time (s) 2.2 2.2 2.2 51.9 54.2
Cost 24.8 24.8 24.8 24.8 24.8
3rd Replan
# Eval. 2087 3359 3629 303122 303122
# Expan. 99080 97759 97131 4100 7989
Time (s) 29.6 27.1 27.4 50.1 52.6
Cost 55.2 55.2 55.2 55.2 55.2
Total Time (s) 461.8 459.3 459.6 664.7 677.1
V Conclusions
@@
We introduce LLPT, an asymptotically optimal lazy lifelong sampling-based path planning algorithm. LLPT combines the replanning efficiency of lifelong planning, the search efficiency of heuristic-guided informed algorithms, and the evaluation efficiency of lazy search algorithms. Our simulations show that LLPT outperforms several sampling-based algorithms for both static and dynamic motion planning problems. These results suggest LLPT is well-suited for time-constrained scenarios, particularly for robots in dynamic environments.

### v3

@@ V Simulation and Experiment Results
In this section, we present simulation results comparing the performance of LLPT vs. other sampling-based planning algorithms with or without lazy search in solving static or dynamic path planning problems  [^1] All simulations were run on Ubuntu 22.04 on an Intel® Core™ Ultra processor (Series 1) CPU machine with 32 GB of RAM.. At the end of this section, we show the results of the proposed algorithm in real-world experiments.
V-A Baseline Algorithms
We compare LLPT against RRT[6], RRT with lazy collision check (denoted as LazyRRT), LazyPRM[47], BIT[49], RRTX[36], DRRT[30], LazyLBTRRT[48], and FAT[37]. In all the plots, We use ’static’ as a suffix to distinguish a lifelong planner from its static version  [^2] The static version can be regarded as an instance of the lifelong version with infinite replanning intervals.. To ensure a fair comparison, each planning algorithm was tested using the Open Motion Planning Library (OMPL) v1.6.0[50]. For the benchmarking algorithms, we used the default OMPL settings. We ensured that LLPT used the same tuning parameters and configurations as the benchmarking algorithms whenever possible. The Euclidean norm is used as the weight and the heuristic functions for all problems. The edge resolution is set as 0.01 for collision check. We use the r-disk RRG as the underlying search graph for all RRT-like planners. The evaluation parameter $\alpha$ of LLPT is set as 100 for all cases.
(a) $\mathbb{SE}(3)$ Cubicles, Scene 1
(b) $\mathbb{SE}(3)$ Cubicles, Scene 2
(c) $\mathbb{SE}(3)$ Cubicles, Scene 3
(d) $\mathbb{SE}(3)$ Piano Movers, Scene 1
(e) $\mathbb{SE}(3)$ Piano Movers, Scene 2. The lamp and the sofa have been moved.
(f) $\mathbb{SE}(3)$ Piano Movers, Scene 3. The wooden desk has been moved.
(g) $\mathbb{R}^{7}$ Manipulation, Scene 1
(h) $\mathbb{R}^{7}$ Manipulation, Scene 2
(i) $\mathbb{R}^{7}$ Manipulation, Scene 3
Figure 5: Depictions of rigid-body motion planning problems in dynamic environments. The dynamic obstacles are highlighted in orange. The shortest collision-free paths for different scenes are displayed in the figures.
V-B Static Motion Planning Problems
We considered the following problems from the OMPL’s test suite:
  • the cubicles problem in $\mathbb{SE}(3)$ as shown in Fig.3(a);
  • the piano movers’ problem in $\mathbb{SE}(3)$ as shown in Fig.3(c).
Besides, we also consider the following problem to investigate the planners’ performance in high-dimensional configuration spaces:
  • the 7-DoF manipulation planning problem on the Franka Emika Panda robot as shown in Fig.3(d), where the Franka Emika Panda robot simulates picking/placing objects between the layers of a bookshelf.
The simulation results are displayed in Fig.4, showcasing success rate and medium solution cost plotted against run time. Several motion planners are not shown due to their performance. At the beginning of the planning, the success rate of LLPT is lower than that of some other planners since it uses more computational time to find an initial solution. However, as planning proceeds, LLPT soon emerges as the top performer among all algorithms in all four scenarios, excelling in success rate and convergence rate towards the optimal solution. These results demonstrate the efficacy of lazily delaying edge evaluations in solving high-dimensional optimal motion planning problems.
V-C Dynamic Motion Planning Problems
The planners were evaluated in three dynamic scenarios where both the positions of obstacles and the robot evolve over time, as illustrated in Fig.5. For each motion planner, we conducted three consecutive planning sessions: the first in Scene 1, the second in Scene 2, and the third in Scene 3. Planners without lifelong planning capabilities must replan from scratch each time the environment changes. TableI reports each planner’s performance in terms of success rate and average solution cost under different replanning time constraints. All results are averaged over 50 trials with different random seeds. To penalize failure cases, the solution cost for unsuccessful trials is set to a large value significantly higher than that of any successful one.
The results demonstrate that LLPT consistently enables effective replanning across all dynamic environments. It achieves substantially better average performance than other planners under the same replanning time limits. In particular, as more replanning time is provided, LLPT significantly outperforms others in terms of average solution cost. Notably, the lifelong planners DRRT and RRTX perform considerably worse than LLPT and even some static planners. DRRT’s subpar performance stems from its lack of asymptotic optimality, while RRTX suffers from high computational overhead due to its costly edge evaluation process. As replanning time decreases, most planners—especially non-lifelong ones—experience a dramatic drop in success rate and a spike in solution cost. In contrast, LLPT maintains a high success rate, low average solution cost, and low standard deviation, demonstrating strong robustness and efficiency in highly dynamic environments.
TABLE I: Success rate (%) and Average Solution Cost of Different Motion Planners under Different Replanning Time for Various Problems Shown in Fig.5
Motion Planner Cubicles Piano Mover Manipulation
T=2s T=5s T=10s T=5s T=10s T=20s T=2s T=5s T=10s
Success Rate (%) RRT 100 100 100 26 54 66 0 0 0
BIT 72 100 100 30 90 100 38 72 98
LazyPRM 94 100 100 2 4 6 32 46 70
LazyRRT 100 100 100 28 46 72 0 0 0
LazyLBTRRT 100 100 100 34 60 66 0 0 0
RRTX 0 0 0 0 0 0 0 0 0
DRRT 88 100 100 20 32 36 66 90 94
LLPT 100 100 100 50 90 100 76 88 94
Solution Cost RRT 1803.5±32.1 1738.9±22.2 1691.9±20.1 388.1±39.4 362.1±34.5 355.2±35.7 - - -
BIT 1895.0±236.0 1737.9±22.7 1699.4±22.1 385.6±43.8 340.0±16.3 329.1±6.1 17.1±8.3 11.0±6.3 6.9±1.3
LazyPRM 1766.7±98.5 1706.1±20.7 1686.7±19.62 469.7±40.1 451.5±50.2 433.9±39.0 17.7±7.4 15.1±7.0 10.7±3.8
LazyRRT 1807.7±23.2 1739.7±21.0 1692.6±21.6 387.3±39.9 363.1±31.2 350.7±33.2 - - -
LazyLBTRRT 1797.4±33.0 1765.6±36.9 1742.5±27.7 386.0±48.9 376.8±32.1 360.7±41.4 - - -
RRTX - - - - - - - - -
DRRT 2051.6±193.8 2058.8±249.9 2024.0±161.2 461.2±61.0 446.4±67.3 442.0±70.4 20.5±5.8 19.5±6.0 19.4±5.4
LLPT 1643.7±18.8 1611.4±17.5 1592.5±16.5 373.3±53.7 331.7±21.0 324.0±4.6 11.3±7.5 8.3±4.9 6.7±2.0
V-D Real-world Experiments
We carried out real-world experiments in a dynamic environment with moving pedestrians. We integrated LLPT with the Nav2 framework [53] and deployed the navigation stack on board the TurtleBot 4 platform (see Fig.6). While moving from the start to the goal position, the robot perceives the environment with the lidar sensor, updates a 2-D grid map of the environment, and updates the search tree using LLPT to avoid static obstacles and pedestrians. Then, a new trajectory is generated and passed to the controller for execution. The whole navigation stack is executed in a real-time manner until the robot reaches the goal position. The experiment demonstrates the algorithm’s ability to handle static and dynamic elements in real-time, ensuring efficient and robust navigation. The video is available at https://youtu.be/49H_wD0pA5U.
Figure 6: The TurtleBot 4 platform for real-world experiments. The robot is equipped with on an onboard computer (NUC11 with i5-1135G7 CPU).
We present some snapshots from experiments in Fig.7. We note that LLPT only carries collision checking around moving pedestrians who block the optimal solution, and only a small number of rewirings are performed for each replan despite the drastic changes of the occupancy map. Besides, LLPT keeps densifying the search tree during the entire navigation procedure, and therefore is able to generate a trajectory with a cost close to the optimal one.
Figure 7: Snapshots captured at 3 distinct time points from the real-world experimental process using LLPT. On the top figures, the search tree of LLPT is plotted with colorful lines representing edges with different statuses. The green lines are the evaluated and valid edges. The red lines are the evaluated and invalid edges. The purple lines are the solution edges.
VI Discussions
VI-A The impact of the evaluation parameter $\alpha$
We explore the impact of the evaluation parameter $\alpha$ on LLPT performance. We compared the LLPT algorithm using various values of $\alpha$, which represents the number of edges evaluated per iteration, in solving dynamic motion planning problems described in Section V-C to assess its impact on performance. The maximum number of nodes of the search graph was fixed for different $\alpha$ to guarantee solution consistency. We summarize the average numbers of total edge evaluations and vertex expansions and the average planning time in 8. The number of vertex expansions refers to the number of vertices that have been updated in the ComputeShortestPath procedure.
We observed that larger values of $\alpha$ result in more edge evaluations but fewer vertex expansions. This is because the larger the number of edges the planner evaluates at the same time, the more knowledge it gains about the state space and, therefore, the less rewiring of the search tree. For the cubicle problem, the replanning time remains relatively stable as $\alpha$ varies. In the Piano Mover scenario, planning time initially decreases at $\alpha=5$ due to reduced vertex expansions, but as $\alpha$ increases further, edge evaluations begin to dominate the computation. For the manipulation problem, edge evaluation is more expensive than vertex expansion. Thus, the replanning time increases as $\alpha$ increases.
The simulation results suggest that the trade-off between edge evaluations and vertex expansions is crucial for determining the convergence rate to the optimal solution. For scenarios where edge evaluation is relatively more expensive compared with vertex expansions, such as robot arm manipulation, or motion planning for systems under differential constraints, a small value of $\alpha$, such as 1, is preferred. In the contrast case where the cost for edge evaluations is computationally insensitive, a larger value of $\alpha$, such as $\infty$, is a better choice since it evaluates all edges on the solution and, therefore, will minimize the number of vertex expansions and facilitate the overall replanning progress.
(a) Cubicles
(b) Piano Movers
(c) Manipulation
Figure 8: Simulation results of LLPT with varying evaluation parameter.
VI-B Performance in Highly-dynamic Scenarios
We compare the performance of LLPT with baseline algorithms across various highly dynamic environments populated by numerous continuously moving obstacles (see Fig.9). Planning results are presented at https://youtu.be/49H_wD0pA5U. The results demonstrate that LLPT can replan and compute optimal solutions at high frequency in response to new obstacle configurations and updated robot positions. This allows LLPT to avoid dynamic obstacles in real time, even without access to their future trajectories. During each replanning step, LLPT evaluates and rewires only a small portion of the search tree near the current solution path. This localized update strategy enables significantly higher robot travel speeds without collisions, outperforming baseline methods.
Figure 9: Simulation environments with continuously-moving obstacles.
VI-C Future Extensions: Improve Convergence Rate by Adaptive Risk-free Region Sampling
Many sampling-based planners exploit the structure of the environment or past planning experience and place samples in promising areas to reduce the time and computational cost of unnecessary exploration. Some of them analyze the geometric structure of the state space and generate random samples from the key area, which lies between the goal and the start region [52] or narrow passages which are difficult to cover by uniform sampling [55, 54]. Recently, learning-based biased sampling is widely explored, which generates the optimal sampling-distribution by learning the probabilistic distribution of free state from past planning experience [56, 57, 58, 59, 60, 61, 62].
Like these works, the convergence rate of LLPT will be further improved if combined with the biased sampling strategy. Nonetheless, most existing works target creating biased samples in static environments. In dynamic environments, the distribution of dynamic obstacles is evolving, and therefore, the optimal sampling distribution should take the future trajectory of dynamic obstacles into consideration to reduce the risk of collision. In this way, an interesting research direction is combining LLPT with adaptive heuristic region sampling that reduces both the collision risk and the overall planning time. Furthermore, as LLPT reuses the search tree throughout the entire navigation process, some parts of the tree should be pruned since they no longer lie in the promising area as the environment evolves. Heuristic-based approaches can be developed to identify and prune the useless tree branches to reduce storage and computation costs.
VII Conclusions
@@
We introduce LLPT, an asymptotically optimal lazy lifelong sampling-based path planning algorithm that combines the replanning efficiency of lifelong planning algorithms, the search efficiency of heuristic-guided informed algorithms, and the evaluation efficiency of lazy search algorithms. The simulation and the experiment results demonstrate that LLPT surpasses several sampling-based planning algorithms to solve static and dynamic motion planning problems in varying dimensions, highlighting its potential applications in high-DoF robotics systems operating in highly dynamic environments.